BlueSky Air has the best on-time arrival rate with 80% of its flights arriving on time . A test is conducted by randomly selecting 16 BlueSky Air flights and observing whether they arrive on time.
Finding the exact probability that at least 8 flights of BlueSky Air arrive late can be very tedious. As such we make use of Normal approximation model. What is the approximated probability? (Please carry answers to at least six decimal places in intermediate steps. Give your final answer to the nearest three decimal places).
Probability (as a proportion)

Answer by Boreal(15235)

mean is np=3.2 expected number of late flights
variance is np(1-p)=3.2*0.8=2.56
sd is sqrt (V)=1.6
at least 8 flights arrive late with expected value of 3.2 and sd of 1.6
normal approximation z>(8.5-3.2)/1.6, using continuity correction factor
This is z>3.3125
probability is 0.0005 or 0.001